# 1in1in1in1in

### v9

@@ Abstract
Reinforcement learning (RL) for physical design of silicon chips in a Google 2021 Nature paper stirred controversy due to poorly documented claims that raised eyebrows and drew critical media coverage. The paper withheld critical methodology steps and most inputs needed to reproduce results. Our meta-analysis shows how two separate evaluations filled in the gaps and demonstrated that Google RL lags behind ($i$) human designers, ($ii$) a well-known algorithm (Simulated Annealing), and ($iii$) generally-available commercial software, while being slower; and in a 2023 open research contest, RL methods weren’t in top 5. Crosschecked data indicate that the integrity of the Nature paper is substantially undermined owing to errors in conduct, analysis and reporting. Before publishing, Google rebuffed internal allegations of fraud. We note policy implications and conclusions for chip design.
@@ 1 Introduction
As AI applications demand greater compute power, efficiency may be improved via better chip design. The Nature paper [1] by Google researchers, published in June 2021, was advertised as a chip-design breakthrough using Machine Learning (ML) [3]. It addressed a challenging problem to optimize locations of circuit components on a chip and described applications to five Tensor Processing Unit (TPU) chip blocks, implying that no better methods were available at the time in academia or industry. The paper broadened the claims beyond chip design to suggest that Reinforcement Learning (RL) extends state of the art in combinatorial optimization.
@@
“Extraordinary claims require extraordinary evidence” [10, 9], but [1] showed no results on public test examples (benchmarks [21]) and did not share the TPU chip blocks used. Source code, released 7 months after publication [4] to support [1] after initial controversy [54, 55, 56, 59, 61], was missing key parts necessary to reproduce the methods and results (as explained in [8, 64]). Over a dozen researchers [54, 55, 61, 64] from Google and academia questioned the claims of [1], ran extensive experiments, and raised further concerns [6, 8] about [1].
@@
Confusingly, the then-head of Google Brain, Dr. Zoubin Ghahramani, a Google VP, tweeted on April 7, 2022 [53] “Google stands by this work published in Nature on ML for Chip Design, which has been independently replicated, open-sourced, and used in production at Google,” apparently referring to reproduction by another Google team (Sergio Guadarrama’s), and without specifying what aspects were reproduced. Google engineers updated their open source [4] many times since, filling in some missing pieces but not all [8]. A single chip-design example was added to [4], but results on it were neither sufficient nor clearly supportive of Google’s RL code [8]. Apparently, the only openly claimed independent (of Google) reproduction of techniques in [1] was developed in Fall 2022 by UCSD researchers [8].111 Efforts by Prof. Andrew B. Kahng at UCSD were praised by Dr. Jeff Dean (the most senior author of the Nature paper [1] and then a Google SVP) in his recorded Dec 2, 2022 workshop keynote [58]. Additional disclosures: as UCSD efforts were starting, Prof. Kahng publicly stated [65] that he was Reviewer #3 of [1]. In the 1990s, Prof. Kahng supervised the doctoral dissertation of the author of this meta-analysis on large-scale VLSI placement at UCLA. They reverse-engineered key components missing from [4] and completely reimplemented the Simulated Annealing (SA) baseline [8] absent in [4]. Google released no TPU chip design blocks used in [1] (nor sanitized equivalents), ruling out full external reproduction of results. So, the UCSD team shared [7] their experiments on modern, public chip designs: SA and commercial EDA tools outperformed Google RL code [4].
@@
Given the earlier exodus of ML researchers from Google, the New York Times and Reuters covered this controversy in 2022 [54, 55]. The reporters found that, well before the Nature submission, Google researchers tasked with checking its claims disputed those claims. Two lead authors of [1] complained of persistent allegations of fraud in their research [56]. In 2022 Google fired an internal whistleblower [54, 55] and denied publication approval for a paper [6] written by Google researchers critical of [1]. The whistleblower sued Google for wrongful termination under California whistleblower-protection laws: court documents [59], filed under penalty of perjury, detail allegations of fraud and scientific misconduct related to research in [1]. The lawsuit moved ahead in Aug ‘23 [66, 68, 74]. Within months of the 2022 media investigations and the lawsuit, the two lead authors of [1] and a senior coauthor left Google [57]. Spring 2023 media coverage noted alleged misrepresentations by Google to potential cloud-services customers [61], questioned reproducibility of results in the Nature paper [64], and covered UCSD research trying to settle the dispute [65]. The 2021 Nature News & Views article introducing [1] in the same issue urged replication of results of [1]. Given the obstacles to and the results of replication [8], the author of the article retracted it.
@@
In this work, Section 2 reviews background and the chip-design task solved in [1, 4]. It also introduces secondary sources used [5, 6, 7, 8]. Section 3 lists initial suspicions about [1]. Section 4 shows that many of them were later confirmed. Section 5 checks if [1] improved the State of the Art (SOTA). Section 6 outlines technical responses by authors of [1] to critiques. Section 7 outlines how the work from [1] may be used in practice. Section 8 draws conclusions about [1] and notes policy implications.
@@
Table 1: Secondary sources published by the teams and chip designs for which they report results. The IBM circuits [23] are ICCAD 2004 benchmarks. [8] built 3 designs with 2 semiconductor technologies.
@@ 2 Background
HPWL stands for half-perimeter wirelength, where the perimeter is taken of the bounding box of points $\{(x_{i},y_{i})\}_{i}$ [30, 40, 52]. It is easy to compute and sum over many nets. This sum correlates with total routed wirelength reasonably well. When $(x,y)$ locations are scaled by a factor $\gamma>0$, HPWL also scales by $\gamma$. This makes HPWL optimization scale-invariant and appropriate for all semiconductor technology nodes. [^2] With semiconductor technology scaling, macros may scale differently, but placement algorithms should handle a variety of macro sizes. Algorithms that optimize HPWL extend to more precisely optimize routed wirelength and technology-dependent chip metrics, so HPWL optimization is a precursor  [19, 24, 31, 34, 36, 37, 38, 39, 40, 42, 44, 45, 49]:
  • to test new placement methods; once HPWL results are close to the best known, accurate metrics are pursued; or
@@
Commercial and academic software for placement is developed to run on modest hardware within reasonable runtime. The methods and software in [1] consume significantly greater resources, but at least with Simulated Annealing (during comparisons) it is straightforward to obtain progressively better results with greater runtime budget.
@@ 3 Initial doubts
While the Nature paper [1] was sophisticated and impressive, its research plan had notable shortfalls. For one, proposed reinforcement learning (RL) was presented as being capable of broader combinatorial optimization (a field that includes puzzle-like tasks such as the Traveling Salesperson Problem, Vertex Cover, Bin Packing). But instead of illustrating this with key problem formulations and easy-to-configure test examples [48], it solved a specialty task (macro placement for chip design) for proprietary Google TPU circuit design blocks, giving results on 5 blocks out of many more available. The RL formulation did not track chip metrics and optimized a simplified proxy function that included HPWL (see Section 2), but was not evaluated for pure HPWL optimization on open circuit examples, as is routine in the literature [19, 21, 24, 25, 26, 27, 31, 30, 36, 40, 45, 49]. New ideas in placement are usually evaluated in research contests on industry chip designs released as public benchmarks [49, Section 6.1], e.g., [30, 37, 38, 39, 40, 42]. but [1] neglected these contest benchmarks.
@@
Figure 1: Layouts from [6, Figure 2] with macros in red and standard cells in green, locations produced by RL (left) and RePlAce (right) for the ibm10 benchmark from [23]. Limiting macro locations to a coarse grid (left) leads to spreading of small macros (red squares on a grid) and elongates connecting wires: from 27.5 (right) to 44.1 (left) for ibm10 [6, Table 1]). High area utilization and many macros of different sizes make the ICCAD 2004 benchmarks [23] challenging compared to [1] and [2, page 43].
@@ 3.2 A flawed optimization proxy
The chip design methodology in [1] uses physical synthesis to generate circuits for further layout optimization (physical design). The proposed RL technique places macros of those circuits to optimize a simplified proxy cost function. Then, a commercial EDA tool is invoked to place the remaining components (standard cells). The remaining operations (including power-grid design, clock-tree synthesis and timing closure [31, 52]) are outsourced to an unknown third party [1, 2]. Results are evaluated with respect to routed wirelength, area, power, and two circuit-timing metrics: TNS and WNS. [^3] TNS = Total Negative Slack, WNS = Worst Negative Slack. These metrics measure violations of timing constraints (negative slack represents violations) by adding violations along all critical paths or using the worst violations. These metrics are noisy since chip timing is often determined by a handful of paths, and small changes to macro locations may change timing a lot.
Per [1], the proxy cost function did not perform circuit timing analysis [52] needed to evaluate TNS and WNS. [^4] Proxy values correlate poorly with TNS and WNS [8]. Therefore, it was misleading to claim in [1] that the proposed RL method led to TNS and WNS improvements on five TPU design blocks without performing variance-based statistical significance tests (TNS and WNS were optimized at later steps unrelated to RL [1]).
@@ 3.3 Use of handicapped techniques
H1. Proposed RL used exorbitant CPU/GPU resources compared to SOTA. Hence, claimed “fast chip design” (presumably, due to fewer unsuccessful design attempts) required careful substantiation.
@@ 3.4 Questionable baselines
B4. [1] did not describe how macro locations in SA were initialized, suggesting that [1] used a naive approach that could be improved. Later, [6] identified more handicaps in the SA baseline in [1], and [8] confirmed them (Section 4).
@@ 4.1 Methods
Undisclosed use of $(x,y)$ locations from commercial tools. Strong evidence and confirmation by Google engineers are mentioned in the UCSD paper [8] that [1] withheld a critical detail. When clustering the input netlist, CT merge in [4] read in a placement to restructure clusters based on locations. To produce $(x,y)$ locations of macros, [1] used initial $(x,y)$ locations of all circuit components (including macros!) produced by commercial EDA tools from Synopsys [4, Issue #25]. The lead authors of [1] confirmed using this step, claiming it was unimportant [60]. But it improved key metrics by 7-10% in [8]. So, the results in [1] needed algorithmic steps absent in [1], e.g., $(x,y)$ data from commercial software.
More undocumented techniques were itemized in [8], which mentioned discrepancies between the Nature paper [1], their source code [4] and the actual code used for chip design at Google. These discrepancies included specific weights of terms in the proxy cost function, a different construction of the adjacency matrix from the circuit, and several “blackbox” elements of [4] available as binaries with no source code or full description in [1]. [6] and [7, 8] offer missing descriptions. Moreover, the results in [1] did not match the methods in [1] because key components were missing. And neither results nor methods were reproducible from descriptions.
Data leakage between training and test data? Per [1], “as we expose the policy network to a greater variety of chip designs, it becomes less prone to overfitting.” But Google Team 1 showed later in [5, Figure 7] that pre-training on “diverse TPU blocks” did not improve quality of results. Pre-training on “previous netlist versions” improved quality somewhat. Pre-training RL and evaluating it on similar designs could be a serious flaw in methodology of [1]. [^6] Such a methodology could help chip designers iterate on design changes faster, but that was not described in [1]. As Google did not release TPU designs or per-design statistics, we cannot compare training and test data.
@@ 5 Did [1] improve SOTA?
The Nature editorial [3] discussing [1] speculated that “this is an important achievement and will be a huge help in speeding up the supply chain.” But today, after evaluations and reproduction attempts at multiple chip-design and EDA companies, it is safe to conclude that no important achievement occurred because prior chip-design software, particularly from Cadence, produced better layouts faster [7, 8]. If this were known to the reviewers of [1] or to the public, the claims of improving TPU designs in [1] would be nonsensical. [1] claimed that humans produced better results than commercial EDA tools, but gave no substantiation. When license terms complicate publishing comparisons to commercial EDA tools, [^7] The lawsuit [59, 66] alleges that Google obtained better results with commercial tools before Nature submission. one compares to academic software and to other prior methods, with the proviso that small improvements are not compelling.
@@ 5.1 Reproduction attempts
[6] and [7, 8] took different approaches to comparing methods from [1] to baselines, but cumulatively reported comparisons to commercial EDA tools, to human designers, to prior university software, and to two independent custom implementations of Simulated Annealing.
@@
  • [6] followed the descriptions in [1] and did not supply initial placement information. [7, 8] sought to replicate what Google actually used to produce results (without description in [1]).
@@
  • Google Team 2 had access to TPU design blocks and evaluated the impact of pre-training in [6]. The impact was small at best. [^8] A consistent conclusion was reported in [5, Figure 7] by Google Team 1 — training on diverse designs does not help, and improvements are seen only when training on earlier versions of the same design. In December 2022 Dr. Jeff Dean, a then Google SVP and the most senior author of [1], also confirmed that RL did well without pre-training [58].
@@
  • The authors of [7, 8] lacked access to Google training data and code but followed Google instructions in [4] for obtaining results similar to those in [1] without pre-training. They also reimplemented SA following instructions in [6] and introduced several new chip-design examples (Table 1).
@@
  • [7, 8] but not [6] performed comparisons using chip metrics and using a commercial EDA tool (Cadence CMP), which outperfomed Google RL. When running RePlace in this context, [8] used only macro locations produced by RePlAce and placed standard cells with the same commercial software used after Google CT/RL [1, 4] (more details below).
@@
  • [7, 8] repeated SA vs. RL comparisons for several configurations (those in [1], those in [4], and additional ones suggested by Google engineers). The results were consistent.
@@
  • [7, 8] demonstrated that a chip designer from IBM outperformed Google RL [4], whereas [6] did not use human baselines.
@@
For comparisons that can be crosschecked, [6] and [7, 8] report qualitatively similar conclusions.
@@
As pointed out in [6], RePlAce was used in [1] in a way inconsistent with its intended use. As a mixed-size placer, RePlAce expects a circuit netlist with macros and standard cells, as a large number of separate components. Instead, the comparisons in [1] suppressed the advantage of RePlAce by clustering standard cells into a few large chunks. [^9] Per Section 2, analytical placers like RePlAce [45] beat other methods on circuits with millions of components. With $<$100K components, earlier methods can be competitive. With proper use of RePlAce, [6] and, independently, [8] produce strong results for RePlAce on well-known public ICCAD 2004 benchmarks. [^10] Comparing HPWL results to those in [25, 27], Google CT/RL [4] underperforms Feng Shui [26] circa 2005. In comparisons on recent designs [7, 8] in terms of chip metrics, RePlAce also shows stronger results, but still loses to Google CT/RL because its placements are harder to route (the losses are much smaller than those reported in [1]). Notably, [6, 7, 8] used RePlAce in a fast mode and not high-quality mode. Also, congestion mitigation was not used, even though RePlace has a congestion-driven mode [45]. In contrast, [1] used routability-improvement techniques, such as cell bloating, without disclosing them (according to [8]). Such techniques can be combined with RePlAce to ensure fair comparisons. Other techniques [24, 31, 36, 40] can be used too.
@@
As explained in [6], the implementation of Simulated Annealing used in [1] was handicapped. Removing the handicaps (in the same source code base) improved results. When properly implemented, SA produces better solutions than Google CT/RL [4] using less runtime, when both are given the same proxy cost function. This is shown consistently in [6, 8] on 17 widely used ICCAD 2004 benchmarks [23] and in [8] on several modern design benchmarks. Compared to Google CT/RL [4], SA consistently improves wirelength and power metrics. For circuit timing metrics TNS and WNS, SA produces less noisy results but comparable to RL’s results [8]. Recall that the proxy function optimized by SA and RL does not include timing metrics [1], making any claims of improvement in these metrics due to SA or RL dubious.

### v10

# 0.70in0.70in0.70in0.75in

@@ Abstract
Reinforcement learning (RL) for physical design of silicon chips in a Google 2021 Nature paper stirred controversy due to poorly documented claims that raised eyebrows and drew critical media coverage. The paper withheld critical methodology steps along with most inputs needed to reproduce results. Our meta-analysis shows how two separate evaluations filled in the gaps and demonstrated that Google RL lags behind ($i$) human chip designers, ($ii$) a well-known algorithm (Simulated Annealing), and ($iii$) generally-available commercial software, while being slower; and in a 2023 open research contest, RL methods weren’t in top 5. Crosschecked data indicate that the integrity of the Nature paper is substantially undermined owing to errors in conduct, analysis and reporting. Before publishing, Google rebuffed internal allegations of fraud. We note policy implications and conclusions for chip design.
@@ 1 Introduction
As AI applications demand greater compute power, efficiency may be improved via better chip design. The Nature paper [1] by Google researchers, published in June 2021, was advertised as a chip-design breakthrough usin Machine Learning (ML) [3]. It addressed a challenging problem to optimize locations of circuit components on a chip and described applications to five Tensor Processing Unit (TPU) chip blocks, implying that no better methods were available at the time in academia or industry. The paper broadened the claims beyond chip design to suggest that Reinforcement Learning (RL) extends state of the art in combinatorial optimization [46].
@@
“Extraordinary claims require extraordinary evidence” [10, 9], but [1] showed no results on public test examples (benchmarks [21]) and did not share the proprietary TPU chip blocks used. Source code, released 7 months after publication [4] to support [1] after initial controversy [55, 56, 57, 61, 63], was missing key parts needed to reproduce the methods and results (as explained in [8, 66]). Over a dozen researchers [55, 56, 63, 66] from Google and academia questioned the claims of [1], performed experiments and raised concerns [6, 8] about [1].
@@
Confusingly, the then-head of Google Brain, Dr. Zoubin Ghahramani, a Google VP, tweeted on April 7, 2022 [54] “Google stands by this work published in Nature on ML for Chip Design, which has been independently replicated, open-sourced, and used in production at Google,” apparently referring to reproduction by another Google team (Sergio Guadarrama’s), and without specifying what aspects were reproduced. Google engineers updated their open source [4] many times since, filling in some missing pieces but not all [8]. A single open-source chip-design example was added to [4], but results on it were neither sufficient nor clearly supportive of Google’s RL code [8]. Apparently, the only openly claimed independent (of Google) reproduction of techniques in [1] was developed in Fall 2022 by UCSD researchers [8].111 Efforts by Prof. Andrew B. Kahng at UCSD were praised by Dr. Jeff Dean (the most senior author of the Nature paper [1] and then a Google SVP) in his recorded Dec 2, 2022 workshop keynote [59]. When UCSD efforts were starting, Prof. Kahng publicly stated [67] that he was Reviewer #3 of [1]. In the 1990s, Prof. Kahng supervised the doctoral dissertation of the author of this meta-analysis on large-scale VLSI placement at UCLA. They reverse-engineered key components missing from [4] and completely reimplemented the Simulated Annealing (SA) baseline [8] absent in [4]. Google released no proprietary TPU chip design blocks used in [1] (nor sanitized equivalents), ruling out full external reproduction of results. So, the UCSD team shared [7] their experiments on modern public chip designs: SA and commercial EDA tools outperformed the Google RL code [4].
@@
Reporters from the New York Times and Reuters covered this controversy in 2022 [55, 56] and found that, well before the Nature submission, several Google researchers disputed the claims they had been tasked with checking. Two lead authors of [1] complained of persistent allegations of fraud in their research [57]. In 2022 Google fired an internal whistleblower [55, 56] and denied publication approval for a paper [6] written by Google researchers critical of [1]. The whistleblower sued Google for wrongful termination under California whistleblower-protection laws: court documents [61], filed under penalty of perjury, detail allegations of fraud and scientific misconduct related to research in [1]. The lawsuit moved ahead in Aug ‘23 [68, 70, 76]. Within months of the 2022 media investigations and the lawsuit, the two lead authors of [1] and a senior coauthor left Google [58]. Spring 2023 media coverage noted alleged misrepresentations by Google to potential cloud-services customers [63], questioned reproducibility of results in the Nature paper [66], and covered UCSD research trying to settle the dispute [67]. The 2021 Nature News & Views article introducing [1] in the same issue urged replication of results of [1]. Given the obstacles to and the results of replication attempts [8], the author of the article retracted it.
@@
In this work, Section 2 reviews background and the chip-design task solved in [1, 4], then introduces secondary sources used [5, 6, 7, 8]. Section 3 lists initial suspicions about [1]. Section 4 shows that many of them were confirmed later. Section 5 checks if [1] improved the State of the Art (SOTA). Section 6 outlines how authors of [1] responded to critiques. Section 7 discusses possible uses of the work in [1] in practice. Section 8 draws conclusions and notes policy implications.
@@
Table 1: Secondary sources published by the teams and chip designs for which they report results. The IBM circuits [23] are ICCAD 2004 benchmarks. [8] built 3 designs with 2 semiconductor technologies each.
@@ 2 Background
HPWL stands for half-perimeter wirelength, where the perimeter is taken of the bounding box of points $\{(x_{i},y_{i})\}_{i}$ [30, 40, 53]. It is easy to compute and sum over many nets. This sum correlates with total routed wirelength reasonably well. When $(x,y)$ locations are scaled by a factor $\gamma>0$, HPWL also scales by $\gamma$. This makes HPWL optimization scale-invariant and appropriate for all semiconductor technology nodes. [^2] With semiconductor technology scaling, some macros may scale differently from logic circuits, but placement algorithms should handle a variety of macro sizes. Algorithms that optimize HPWL extend to more precisely optimize routed wirelength and technology-dependent chip metrics, so HPWL optimization is a precursor  [19, 24, 31, 34, 36, 37, 38, 39, 40, 42, 44, 45, 50]:
  • to test new placement methods; once HPWL results are close to the best known, more accurate metrics are used for evaluation; or
@@
Commercial and academic software for placement is developed to run on modest hardware within reasonable runtime. The methods and software in [1] consume much greater resources, but at least with Simulated Annealing (during comparisons) it is straightforward to obtain progressively better results with greater runtime budget.
@@ 3 Initial doubts
While the Nature paper [1] was sophisticated and impressive, its research plan had notable shortfalls. For one, the proposed reinforcement learning (RL) was presented as being capable of broader combinatorial optimization (a field that includes puzzle-like tasks such as the Traveling Salesperson Problem, Vertex Cover, Bin Packing). But instead of illustrating this with key problem formulations and easy-to-configure test examples [49], it solved a specialty task (macro placement for chip design) for proprietary Google TPU circuit design blocks, giving results on 5 blocks out of many more available. The RL formulation did not track chip metrics and optimized a simplified proxy function that included HPWL (see Section 2), but was not evaluated for pure HPWL optimization on open circuit examples, as is routine in the literature [19, 21, 24, 25, 26, 27, 31, 30, 36, 40, 45, 50]. New ideas in placement are usually evaluated in research contests on industry chip designs released as public benchmarks [50, Section 6.1], e.g., [30, 37, 38, 39, 40, 42]. but [1] neglected these contest benchmarks.
@@
Figure 1: Layouts from [6, Figure 2] with macros in red and standard cells in green, locations produced by RL (left) and RePlAce (right) for the ibm10 benchmark from [23]. Limiting macro locations to a coarse grid (left) leads to spreading of small macros (red squares on the grid) and elongates connecting wires: from 27.5 units (right) to 44.1 units (left) for ibm10 [6, Table 1]). Higher area utilization and many macros of different sizes make the ICCAD 2004 benchmarks [23] challenging compared to benchmarks in [1] and [2, page 43].
@@ 3.2 A flawed optimization proxy
The chip design methodology in [1] uses physical synthesis to generate circuits for further layout optimization (physical design). The proposed RL technique places macros of those circuits to optimize a simplified proxy cost function. Then, a commercial EDA tool is invoked to place the remaining components (standard cells). The remaining operations (including power-grid design, clock-tree synthesis and timing closure [31, 53]) are outsourced to an unknown third party [1, 2]. Results are evaluated with respect to routed wirelength, area, power, and two circuit-timing metrics: TNS and WNS. [^3] TNS = Total Negative Slack, WNS = Worst Negative Slack. These metrics measure violations of timing constraints (negative slack represents violations) by adding violations along all critical paths or using the worst violations. These metrics are noisy since chip timing is often determined by a handful of paths, and small changes to macro locations may change timing a lot. Per [1], the proxy cost function did not perform circuit timing analysis [53] needed to evaluate TNS and WNS. [^4] Proxy values correlate poorly with TNS and WNS [8]. Therefore, it was misleading to claim in [1] that the proposed RL method led to TNS and WNS improvements on five TPU design blocks without performing variance-based statistical significance tests (TNS and WNS were optimized at later steps unrelated to RL [1]).
@@ 3.3 Use of handicapped techniques
H1. The proposed RL used exorbitant CPU/GPU resources compared to SOTA. Hence, claimed “fast chip design” (presumably, due to fewer unsuccessful design attempts) required careful substantiation.
@@ 3.4 Questionable baselines
B4. [1] did not describe how macro locations in SA were initialized, hinting that the naive approach in [1] could be improved. Later, [6] identified more handicaps in the SA baseline in [1], and [8] confirmed them (Section 4).
@@ 4.1 Methods
Undisclosed use of $(x,y)$ locations from commercial tools. Strong evidence and confirmation by Google engineers are mentioned in the UCSD paper [8] that [1] withheld a critical detail. When clustering the input netlist, CT merge code in [4] read in a placement to restructure clusters based on locations. To produce $(x,y)$ locations of macros, [1] used initial $(x,y)$ locations of all circuit components (including macros!) produced by commercial EDA tools from Synopsys [4, Issue #25]. The lead authors of [1] confirmed using this step, claiming it was unimportant [62]. But it improved key metrics by 7-10% in [8]. So, the results in [1] needed algorithmic steps absent in [1], e.g., obtaining $(x,y)$ data from commercial software.
More undocumented techniques were itemized in [8], which mentioned discrepancies between the Nature paper [1], their source code [4] and the actual code used for chip design at Google. These discrepancies included specific weights of terms in the proxy cost function, a different construction of the adjacency matrix from the circuit, and several “blackbox” elements of [4] available as binaries with no source code or full description in [1]. [6] and [7, 8] offer missing descriptions. Moreover, the results in [1] did not match the methods in [1] because key components were missing. And neither results nor methods were reproducible from descriptions alone.
Data leakage between training and test data? Per [1], “as we expose the policy network to a greater variety of chip designs, it becomes less prone to overfitting.” But Google Team 1 showed later in [5, Figure 7] that pre-training on “diverse TPU blocks” did not improve quality of results. Pre-training on “previous netlist versions” improved quality somewhat. Pre-training RL and evaluating it on similar designs [80] might be a serious flaw in methodology of [1]. [^6] Such a methodology could help chip designers iterate on design changes faster, but that was not described in [1]. As Google did not release proprietary TPU designs or per-design statistics, we cannot compare training and test data.
@@ 5 Did [1] improve SOTA?
The Nature editorial [3] discussing [1] speculated that “this is an important achievement and will be a huge help in speeding up the supply chain.” But today, after evaluations and reproduction attempts at multiple chip-design and EDA companies, it is safe to conclude that no important achievement occurred because prior chip-design software, particularly from Cadence Design Systems, produced better layouts faster [7, 8]. If this were known to the reviewers of [1] or to the public, the claims of improving TPU designs in [1] would be nonsensical. [1] claimed that humans produced better results than commercial EDA tools, but gave no substantiation. When license terms complicate publishing comparisons to commercial EDA tools, [^7] The lawsuit [61, 68] alleges that Google obtained better results with commercial tools before Nature submission. one compares to academic software and to other prior methods, with the proviso that small improvements are not compelling.
@@ 5.1 Reproduction attempts
Google Team 2 [6] and the UCSD Team [7, 8] took different approaches to comparing methods from [1] to baselines, but cumulatively reported comparisons to commercial EDA tools, to human designers, to prior university software, and to two independent custom implementations of Simulated Annealing.
@@
  • Google Team 2 in [6] followed the descriptions in [1] and did not supply initial placement information. The UCSD Team in [7, 8] sought to replicate what Google actually used to produce results (without description in [1]).
@@
  • Google Team 2 had access to TPU design blocks and evaluated the impact of pre-training in [6]. The impact was small at best. [^8] A consistent conclusion was reported in [5, Figure 7] by Google Team 1 — training on diverse designs does not improve quality of results, and improvements are seen only when training on earlier versions of the same design. In December 2022 Dr. Jeff Dean, a then Google SVP and the most senior author of [1], also confirmed that RL did well without pre-training [59].
@@
  • The UCSD Team [7, 8] lacked access to Google training data and code but followed Google instructions in [4] for obtaining results similar to those in [1] without pre-training. They also reimplemented SA following instructions in [6] and introduced several new chip-design examples (Table 1).
@@
  • The UCSD Team [7, 8] but not Google Team 2 [6] performed comparisons using chip metrics and using a commercial EDA tool (Cadence CMP), where Cadence CMP outperfomed Google RL. When running RePlAce in this context, [8] used only macro locations produced by RePlAce and placed standard cells with the same commercial software used after Google CT/RL [1, 4] (more details below).
@@
  • The UCSD Team [7, 8] repeated SA vs. RL comparisons for several configurations (those in [1], those in [4], and additional ones suggested by Google engineers). The results were consistent.
@@
  • The UCSD Team [7, 8] demonstrated that a chip designer from IBM outperformed Google RL [4], whereas Google Team 2 [6] did not use human baselines.
@@
For comparisons that can be crosschecked, the two teams report (in [6] and [7, 8]) qualitatively similar conclusions.
@@
As pointed out in [6], RePlAce was used in [1] in a way inconsistent with its intended use. As a mixed-size placer, RePlAce expects a circuit netlist with macros and standard cells, as a large number of separate components. Instead, the comparisons in [1] suppressed the advantage of RePlAce by clustering standard cells into a few large chunks. [^9] Per Section 2, analytical placers like RePlAce [45] beat other methods on circuits with millions of components. With $<$100K components, earlier methods can be competitive. With proper use of RePlAce, [6] and, independently, [8] obtain strong results with RePlAce on well-known public ICCAD 2004 benchmarks. [^10] Comparing HPWL results to those in [25, 27], Google CT/RL [4] underperforms Feng Shui [26] circa 2005. In comparisons on recent designs [7, 8] RePlAce also attains better chip metrics, but still loses to Google CT/RL because its placements are harder to route (the losses are much smaller than those reported in [1]). Notably, [6, 7, 8] used RePlAce in a fast mode and not high-quality mode. Congestion-driven mode in RePlAce [45] was not used. In contrast, [1] used routability-improvement techniques, such as cell bloating, without disclosing them (according to [8]). Such techniques can be combined with RePlAce to ensure fair comparisons. Other techniques [24, 31, 36, 40] can be used too.
@@
As explained in [6], the implementation of Simulated Annealing used in [1] was handicapped. Removing the handicaps (in the same source code base) improved results. When properly implemented, SA produces better solutions than Google CT/RL [4] using less runtime, when both are given the same proxy cost function. This is shown consistently in [6, 8] on 17 widely used ICCAD 2004 benchmarks [23] and in [8] on several modern design benchmarks. Compared to Google CT/RL [4], SA consistently improves wirelength and power metrics. For circuit timing metrics TNS and WNS, SA produces less noisy results that are otherwise comparable to RL’s results [8]. Recall that the proxy function optimized by SA and RL does not include timing metrics [1], making any claims of improvement in these metrics due to SA or RL dubious.